OpenStudy (anonymous):

Given two functions f and g, one way to obtain the real solutions of the equation f(x)=g(x) is to graph the equation y=f(x) and y=g(x) in an xy-plane and then find the x-coordinates of any points of intersection. Describe another way to solve f(x)=g(x) that also involves graphing in an xy-plane but that is based on the difference function f-g. THANKS SO MUCH

OpenStudy (anonymous):

let h(x) = f(x) - g(x) since f(x) = g(x), you'd be finding the roots of h(x). In other words, you'd be solving h(x) = 0. This equivalent to finding where h(x) crosses the x axis.

OpenStudy (anonymous):

thanks a ton!

OpenStudy (anonymous):

yw...:)

OpenStudy (anonymous):

how would you solve \[\sqrt{x+1} = 2x\] using that though?